OpenStudy (anonymous):

find the area of a triangle with the base of 1and 2 thirds yards and a height 3 sevenths yard

OpenStudy (anonymous):

find the area of a triangle with the base of \[1\frac{ 2 }{ 3 }\] yards and a height\[\frac{ 3 }{ 7 }\]yard

OpenStudy (anonymous):

area of triangle = (1/2)xbasexheight base = ((1x3)+2)/3=5/3 height = 3/7 area= (1/2)x(5/3)x(3/7) = 15/42=5/14

OpenStudy (anonymous):

yard^2

OpenStudy (mayankdevnani):

area of triangle=1/2*b*h

OpenStudy (mayankdevnani):

ok

OpenStudy (anonymous):

thanks that helps me understand it

OpenStudy (mayankdevnani):

ok